# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.1.2 Item 3: Subsampling from the block source
We denote the corresponding samples by $W_{i}=\Gamma_{i}(Y|_{[1,d_{i}]})$, and let $S_{i}=Z_{i}|_{W_{i}}$. Setting $\varepsilon_{i}^{\prime}=2^{-(\zeta/2)\delta\Delta_{i}}$ and observing that $\delta\Delta_{i}=(1-\frac{\zeta}{2})\delta\Delta_{i}+\log(1/\varepsilon^{\prime}_{i})$, we get that $Z$ is $\varepsilon^{\prime}=\sum_{i}\varepsilon^{\prime}_{i}$ close to some $Z^{\prime}$, an exact $((\Delta_{1},\ldots,\Delta_{t}),(1-\zeta)\delta)$-source. From here onwards, assume that $Z$ is the exact block source, and aggregate the error.
Next, we invoke Lemma 2.16 with $\tau=\frac{\zeta\delta}{6}$ (notice that indeed $\varepsilon_{\Gamma}\leq\frac{\tau}{\log(1/\tau)}$), and get that for every $i\in[t]$, and $z_{\mathsf{pre}}\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\Delta_{1}+\ldots+\Delta_{i-1}}$,
$S_{i}\mid\mathopen{}\mathclose{{}\left\{{(Z_{1},\ldots,Z_{i-1})=z_{\mathsf{pre}}}}\right\}$
is $\varepsilon^{\prime\prime}_{i}=\gamma+2^{-\Omega(\tau\Delta_{i})}$-close to having min-entropy $(1-\frac{\zeta}{2})^{2}\delta\cdot\ell_{i}\geq(1-\zeta)\delta\cdot\ell_{i}$. Thus, in particular, it holds if we condition on any sample from $(S_{1},\ldots,S_{i-1})$, and so we have that for every $i\in[t]$,
$\mathopen{}\mathclose{{}\left(S_{1},\ldots,S_{i-1},S_{i}}\right)\approx_{\varepsilon^{\prime\prime}_{i}}\mathopen{}\mathclose{{}\left(S_{1},\ldots,S_{i-1},S^{\prime}_{i}}\right),$
where $S^{\prime}_{i}$ has $(1-\zeta)\delta$ entropy rate. This means [^17] In what follows, we use the fact that we can couple any two $X\approx_{\varepsilon}X^{\prime}$ with $(X,Y)\approx_{\varepsilon}(X^{\prime},Y)$, for any joint distribution $(X,Y)$. See, e.g., [Li15, Lemma 3.20]. that $(S_{1},\ldots,S_{t})$ has distance
@@
from an $((\ell_{1},\ldots,\ell_{t}),(1-\zeta)\delta)$ block source, where we used the fact that the $2^{-\Omega(\tau\Delta_{1})}$ and $2^{-(\zeta/2)\delta\Delta_{1}}$ terms are at most $\frac{\varepsilon}{4t}$, which follows from the fact that $c_{1}\log(t/\varepsilon)\leq\Delta_{1}$ for a suitable choice of $c_{1}$.
@@
$\sum_{i=1}^{t}\ell_{i}\cdot\operatorname{polylog}(\Delta_{i})=\operatorname{polylog}(\Delta_{t})\cdot\ell_{1}$
@@
5.1.3 Item 4: Applying a block source extractor
We now wish to extract from our decreasing-blocks block source, and for that we combine Lemmas 5.3 and 5.2 with the block source extraction of Lemma 2.23, which will give us a nearly linear-time logarithmic-seed extractor that outputs $n^{\Omega(1)}$ bits. For the $\mathsf{Ext}_{i}$-s in Lemma 2.23, we will use the fast hash-based extractors from Lemma 2.14.
@@ 5.1.3 Item 4: Applying a block source extractor
There exists a small constant $c>0$ such that the following holds. For every large enough $n$, any constant $\delta\in(0,1)$, any $k\geq\delta n$, and any $\varepsilon\geq 2^{-n^{c}}$, there exists a $(k,\varepsilon)$ extractor
@@
Let $X$ be an $(n,k=\delta n)$-source. Set $\varepsilon^{\prime}=\varepsilon/3$, $\theta_{1}=8/10$, $\theta_{2}=9/10$, and $\zeta=1/10$. We first apply Lemma 5.2 with $\Delta_{t}=n^{\theta_{2}}$, $\Delta_{1}=n^{\theta_{1}}$, and error $\varepsilon^{\prime}$, where $t=O(\log n)$ is as guaranteed from the lemma’s statement. This requires a seed of length $d_{1}=O(\log(1/\varepsilon^{\prime}))=O(\log(1/\varepsilon))$, and in time $\widetilde{O}(n)$ we output a random variable $Z_{1}$ which is $\varepsilon^{\prime}$-close to a $((\Delta_{1},\ldots,\Delta_{t}),(1-\zeta)\delta)$ block source. Assume that $Z_{1}$ is exactly a block source, and aggregate the error.
Set $\beta=7/10$, and $\gamma=6/10<\beta$. Set $\alpha$ to be the constant such that $n^{\beta}\cdot\alpha^{t-1}=n^{\gamma}$. We then apply Lemma 5.3 on $Z_{1}$ with that $\alpha$, the same $\zeta$, closeness $\varepsilon^{\prime}$ and an initial block length $\ell_{1}=n^{\beta}$.This gives us a random variable $Z_{2}$ that is $2\varepsilon^{\prime}$-close to a
@@
block source, requires a seed of length $d_{2}=O(\log(n/\varepsilon^{\prime}))=O(\log(n/\varepsilon))$, and runs in time $O(n)$. Again, assume that $Z_{2}$ is exactly a block source, and aggregate the error.
For our next and final step, set $d_{3}=c_{\mathsf{E}}\log(\ell_{t}/\varepsilon_{\mathsf{Ext}})$ where $c_{\mathsf{E}}$ is the constant guaranteed by Lemma 2.14. Also, let $\varepsilon_{\mathsf{Ext}}=\frac{\varepsilon^{\prime}}{6t}$, and $\theta$ will be a constant whose value will be later determined. We will use the following extractors:
  • Let $\mathsf{Ext}_{t}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell_{t}}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{3}}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m_{t}=(1+\theta)d_{3}}$ be the $(k_{t}=(\delta^{\prime}/2)\ell_{t},\varepsilon_{\mathsf{Ext}})$ extractor guaranteed to us by Lemma 2.14. Notice that we need to satisfy $k_{t}\geq\theta d_{3}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})$. Looking forward, we will also need that $(\delta^{\prime}/2)\ell_{t}\leq\delta^{\prime}\ell_{t}-\log(1/\varepsilon_{\mathsf{Ext}})$. Those constraints can be satisfied making sure that $\varepsilon$ is at most $2^{-\Omega(\ell_{t})}$, where the hidden constant depends on $c_{\mathsf{E}}$.
@@
    be the $(k_{i}=(\delta^{\prime}/2)\ell_{i},\varepsilon_{\mathsf{Ext}})$ extractor guaranteed to us by Lemma 2.14, where $m_{i}=(1+\theta)m_{i+1}$. We need to make sure that $m_{i+1}\geq c_{\mathsf{E}}\log(\ell_{i}/\varepsilon_{\mathsf{Ext}})$ and that $k_{i}\geq\theta m_{i+1}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})$. To see that the latter holds, note that $k_{i}=(\delta^{\prime}/2)\ell_{1}\cdot\alpha^{i-1}\geq n^{\gamma/2}$ and that $\theta m_{i+1}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})=\theta(1+\theta)^{t-i}d_{3}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})<n^{\gamma/2},$ if we choose $\theta$ to be a small enough constant (with respect to the constant $\frac{\log n}{t}$) and $\varepsilon$ to be, again, at most $2^{-\Omega(\ell_{t})}$. Also, here too, record that $(\delta^{\prime}/2)\ell_{i}\leq\delta^{\prime}\ell_{i}-\log(1/\varepsilon_{\mathsf{Ext}})$, which follows easily, since the $\ell_{i}$-s increase.
Everything is in place to apply our block source extraction, Lemma 2.23, on $Z_{2}$ and an independent and uniform seed of length $d_{3}$. [^18] Note that here, we use Lemma 2.23 with $n_{i}=(1+\theta)^{-i}n_{1}$ and $k_{i}\geq\theta n_{i}-\log(1/\varepsilon_{i})$. The slack in entropy is needed since we work with the notion of block sources that also allows average conditional min-entropy. We can thus use the fact that under such setting of parameters, every extractor is an average case extractor with only a slight loss in parameters (see, e.g., [DORS08]). We omit the easy proof. We get that $\mathsf{BSExt}$ outputs $Z_{3}$ of length $m_{1}=n^{\Omega(1)}$, which is $2t\varepsilon_{\mathsf{Ext}}\leq\varepsilon^{\prime}$ close to uniform, and runs in time $O\mathopen{}\mathclose{{}\left(\sum_{i=1}^{t}\ell_{i}\log\ell_{i}}\right)=O(n).$
@@ 5.1.4 Improving the output length
The extractor $\mathsf{Ext}_{\mathsf{short}}$ from Lemma 5.4 only outputs $n^{\Omega(1)}$ bits. Here, we will use an extractor $\mathsf{Ext}_{\mathsf{aux}}$ that outputs a linear fraction of the entropy but requires a (relatively) long seed, and use Lemma 2.26 to boost the output length. For $\mathsf{Ext}_{\mathsf{aux}}$, we will again use a sample-then-extract extractor, however this time, we can use independent samples to create a block source with exponentially decreasing blocks. This setting is easier, and we can simply use the original [NZ96] construction. Since a similar construction will be analyzed later in the paper (including a time complexity analysis), we choose to employ it instead of revisiting [NZ96].
@@
There exist constants $\tau,c\in(0,1)$ and $C>1$, such that for every positive integer $n$, and any $\varepsilon\geq 2^{-n^{c}}$, there exists a strong $(k=(1-\tau)n,\varepsilon)$ extractor
@@
where $d=O(\log n\cdot\log(n/\varepsilon))$, and $m=ck-C\log(1/\varepsilon)$. Moreover, given inputs $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$, we can compute $\mathsf{Ext}_{\mathsf{out}}(x,y)$ in time $\widetilde{O}(n)$.
The correctness follows from Lemma 5.9 (without the need for a preliminary condensing step), employed with the hash functions of Lemma 2.14.
Plugging-in $\mathsf{Ext}_{\mathsf{out}}$ and $\mathsf{Ext}_{\mathsf{short}}$ into Lemma 2.26 readily gives the following result.
@@
There exist constants $\tau,c\in(0,1)$ such that for every positive integer $n$, and any $\varepsilon\geq 2^{-n^{c}}$, there exists a $(k=(1-\tau)n,\varepsilon)$ extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ where $d=O(\log(n/\varepsilon))$, and $m=ck$. Moreover, given inputs $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$, we can compute $\mathsf{Ext}(x,y)$ in time $\widetilde{O}(n)$.
To boost the output length from $\Omega(k)$ to $(1-\eta)k$ for any constant $\eta>0$, we apply Lemma 2.22 a constant number of times depending only on $\eta$ (that is, we simply apply $\mathsf{Ext}$ with independent seeds and concatenate the outputs). To go from any min-entropy $k$ to entropy rate $1-\tau$, we first apply a condenser, either the one from Theorem 3.10 or the one from Theorem 3.11. Specifically, when $k\geq C\log^{2}(n/\varepsilon)$, we can use Theorem 3.10 which takes $\widetilde{O}(n)$ time. When $k$ is smaller, we can use Theorem 3.11, but this requires an extra preprocessing time which takes $T_{\mathsf{pre}}=\operatorname{polylog}(1/\varepsilon)$ times. Note that the bound on $\varepsilon$ from Lemma 5.6 translates to $\varepsilon\geq 2^{-k^{c}}$, so we can (if needed) modify $c$ so that $T_{\mathsf{pre}}=O(n)$. This finally gives us our main theorem for this section, Theorem 5.1, apart from the strongness property, which we now discuss.
@@
In what follows, we refer to the itemized list in the beginning of the section. The condensing step, Item 1, is strong, since we use strong condensers. Next, inspecting the proofs of Lemmas 5.2 and 5.3, we see that both samplings procedures yield a good sample with high probability over the fixing of the seed, so Items 2 and 3 hold in a strong manner as well. Item 4 follows by applying a block source extractor, which is strong since the extraction steps output the seed. Thus, the extractor $\mathsf{Ext}_{\mathsf{short}}$ from Lemma 5.4 is in fact strong. For the output-extending phase, Lemma 2.26 readily tells us that the extractor from Lemma 5.6 is strong. Finally, we apply that extractor several times with independent seeds, and the strongness of that procedure is guaranteed from Lemma 2.22.

### v2

@@ 4.1 A Non-Recursive Construction
We denote the corresponding samples by $W_{i}=\Gamma_{i}(Y|_{[1,d_{i}]})$, and let $S_{i}=(Z_{i})_{W_{i}}$. Setting $\varepsilon_{i}^{\prime}=2^{-(\zeta/2)\delta\Delta_{i}}$ and observing that $\delta\Delta_{i}=(1-\frac{\zeta}{2})\delta\Delta_{i}+\log(1/\varepsilon^{\prime}_{i})$, we get that $Z$ is $\varepsilon^{\prime}=\sum_{i}\varepsilon^{\prime}_{i}$ close to some $Z^{\prime}$, an exact $((\Delta_{1},\ldots,\Delta_{t}),(1-\zeta)\delta)$-source. From here onwards, assume that $Z$ is the exact block source, and aggregate the error.
Next, we invoke lemma˜2.15 with $\tau=\frac{\zeta\delta}{6}$ (notice that indeed $\varepsilon_{\Gamma}\leq\frac{\tau}{\log(1/\tau)}$), and get that for every $i\in[t]$, and $z_{\mathsf{pre}}\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\Delta_{1}+\ldots+\Delta_{i-1}}$,
$\mathopen{}\mathclose{{\left(Y,S_{i,z_{\mathsf{pre}}}}}\right)\approx_{\varepsilon^{\prime\prime}_{i}=\gamma+2^{-\Omega(\tau\Delta_{i})}}\mathopen{}\mathclose{{\left(Y,S^{\prime}_{i,z_{\mathsf{pre}}}}}\right),$
where $S_{i,z_{\mathsf{pre}}}$ denotes $S_{i}$ conditioned on $(Z_{1},\dots,Z_{i-1})=z_{\mathsf{pre}}$ and $S^{\prime}_{i,z_{\mathsf{pre}}}$ satisfies $\mathbf{H}_{\infty}(S^{\prime}_{i,z_{\mathsf{pre}}}|Y=y)\geq(1-\frac{\zeta}{2})^{2}\delta\cdot\ell_{i}\geq(1-\zeta)\delta\cdot\ell_{i}$ for all $y$. Thus, in particular, this holds if we condition on any sample from $(S_{1},\ldots,S_{i-1})$, and so we have that for every $i\in[t]$,
$\mathopen{}\mathclose{{\left(Y,S_{1},\ldots,S_{i-1},S_{i}}}\right)\approx_{\varepsilon^{\prime\prime}_{i}}\mathopen{}\mathclose{{\left(Y,S_{1},\ldots,S_{i-1},S^{\prime}_{i}}}\right),$ (7)
where $\mathbf{H}_{\infty}(S^{\prime}_{i}|Y=y,S_{1}=s_{1},\dots,S_{i-1}=s_{i-1})\geq(1-\zeta)\delta\cdot\ell_{i}$ for all $(y,s_{1},\dots,s_{i-1})$.
This implies that, conditioned on the seed $Y$, $(S_{1},\ldots,S_{t})$ has distance
@@
from an (exact) $((\ell_{1},\ldots,\ell_{t}),(1-\zeta)\delta)$ block source, where we used the fact that the $2^{-\Omega(\tau\Delta_{1})}$ and $2^{-(\zeta/2)\delta\Delta_{1}}$ terms are at most $\frac{\varepsilon}{4t}$, which follows from the fact that $c_{1}\log(t/\varepsilon)\leq\Delta_{1}$ for a suitable choice of $c_{1}$, and where $\varepsilon^{\prime}$ accounts for the assumption that $Z$ is the exact block source above. This can be shown by induction on the number of blocks using eq.˜7 analogously to (and even in a simpler way than) [NZ96, proof of Lemma 17], and similarly to the proof of lemma˜4.2. Since we believe this proof is easier to follow than the proof of lemma˜4.2, we give details here too for the sake of exposition.
First, the base case is given by eq.˜7 for $i=1$. Now, fix an arbitrary $i\geq 2$ and suppose that we already know that
$(Y,S_{1},\dots,S_{i-1})\approx_{\sum_{j=1}^{i-1}\varepsilon^{\prime\prime}_{j}}(Y,S^{\prime}_{1},\dots,S^{\prime}_{i-1}),$ (8)
where $\mathbf{H}_{\infty}(S^{\prime}_{j}|Y=y,S^{\prime}_{1}=s_{1},\dots,S^{\prime}_{j-1}=s_{j-1})\geq(1-\zeta)\delta\ell_{j}$ for every $1\leq j\leq i-1$ and all $(y,s_{1},\dots,s_{j-1})$.
We now show how to extend this by one block. Generally speaking, suppose that we already have blocks $B_{1},\dots,B_{i-1}$, arbitrarily distributed. Analogously to the proof of lemma˜4.2, we generate one more block $B_{i}$ by first sampling $\vec{v}\sim(Y,B_{1},\dots,B_{i-1})$. If $\vec{v}$ is in the support of $(Y,S_{1},\dots,S_{i-1})$, we set $B_{i,\vec{v}}$, the random variable $B_{i}$ conditioned on $(Y,B_{1},\dots,B_{i-1})=\vec{v}$, to be $B_{i,\vec{v}}=(S^{\prime}_{i}|(Y,S_{1},\dots,S_{i-1})=\vec{v})$. Otherwise, we set $B_{i,\vec{v}}$ to be uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell_{i}}$ and independent of everything else. By construction, $\mathbf{H}_{\infty}(B_{i,\vec{v}})\geq(1-\zeta)\delta\ell_{i}$ for all $\vec{v}$.
From eq.˜8, it follows that
$\mathopen{}\mathclose{{\left(Y,S_{1},\dots,S_{i-1},B^{(1)}_{i}}}\right)\approx_{\sum_{j=1}^{i-1}\varepsilon^{\prime\prime}_{j}}\mathopen{}\mathclose{{\left(Y,S^{\prime}_{1},\dots,S^{\prime}_{i-1},B^{(2)}_{i}}}\right)$, (9)
where $B^{(1)}_{i}$ is sampled by following the procedure in the paragraph above with $B_{j}=S_{j}$ for all $j\leq i-1$, and $B^{(2)}_{i}$ is sampled from $B_{j}=S^{\prime}_{j}$ for all $j\leq i-1$. Now, note that $(Y,S_{1},\dots,S_{i-1},B^{(1)}_{i})$ is distributed exactly like $(Y,S_{1},\dots,S_{i-1},S^{\prime}_{i})$, because when $B_{j}=S_{j}$ for all $j\leq i-1$ we get that $\vec{v}$ above is always in the correct support. Therefore, combining this observation with eqs.˜7 and 9 and a triangle inequality yields
$(Y,S_{1},\dots,S_{i-1},S_{i})\approx_{\varepsilon^{\prime\prime}_{i}+\sum_{j=1}^{i-1}\varepsilon^{\prime\prime}_{j}=\sum_{j=1}^{i}\varepsilon^{\prime\prime}_{j}}\mathopen{}\mathclose{{\left(Y,S^{\prime}_{1},\dots,S^{\prime}_{i-1},B^{(2)}_{i}}}\right)$.
To conclude the argument, it suffices to note that $(S^{\prime}_{1},\dots,S^{\prime}_{i-1},S^{\prime}_{i}=B^{(2)}_{i}|Y=y)$ is the desired exact block source by inspection of the sampling process for the $S^{\prime}_{j}$-s, and take $i=t$.
@@
$\sum_{i=1}^{t}\log^{2}(1/\gamma)\cdot\operatorname{polylog}(\Delta_{i})+O(\ell_{i}\log\Delta_{i})\leq t\cdot\operatorname{polylog}(\Delta_{t})\mathopen{}\mathclose{{\left(\ell_{1}+\log^{2}(t/\varepsilon)}}\right)$
@@
4.1.3 item˜4: Applying a block source extractor
We now wish to extract from our decreasing-blocks block source, and for that we combine lemmas˜4.3 and 4.2 with the block source extraction of lemma˜2.22. This will give us a nearly linear-time logarithmic-seed extractor that outputs $n^{\Omega(1)}$ bits. For the $\mathsf{Ext}_{i}$-s in lemma˜2.22, we will use the fast hash-based extractors from lemma˜2.13.
@@
There exists a small constant $c>0$ such that the following holds. For every large enough $n$, any constant $\delta\in(0,1)$, any $k\geq\delta n$, and any $\varepsilon\geq 2^{-n^{c}}$, there exists a strong $(k,\varepsilon)$ extractor
@@
Let $X$ be an $(n,k=\delta n)$-source. Set $\varepsilon^{\prime}=\varepsilon/3$, $\theta_{1}=8/10$, $\theta_{2}=9/10$, and $\zeta=1/10$. We first apply lemma˜4.2 with $\Delta_{t}=n^{\theta_{2}}$, $\Delta_{1}=n^{\theta_{1}}$, and error $\varepsilon^{\prime}$, where $t=O(\log n)$ is as guaranteed from lemma˜4.2. This requires a seed of length $d_{1}=O(\log(1/\varepsilon^{\prime}))=O(\log(1/\varepsilon))$, and in time $\widetilde{O}(n)$ we output a random variable $Z_{1}$ which is $\varepsilon^{\prime}$-close to an exact $((\Delta_{1},\ldots,\Delta_{t}),(1-\zeta)\delta)$ block source for every fixing of the seed.
Set $\beta=7/10$, and $\gamma=6/10<\beta$. Set $\alpha$ to be the constant such that $n^{\beta}\cdot\alpha^{t-1}=n^{\gamma}$. We then apply lemma˜4.3 on $Z_{1}$ with that $\alpha$, the same $\zeta$, closeness $\varepsilon^{\prime}$ and an initial block length $\ell_{1}=n^{\beta}$. This gives us a random variable $Z_{2}$ that is $2\varepsilon^{\prime}$-close to a
@@
block source, requires a seed of length $d_{2}=O(\log(n/\varepsilon^{\prime}))=O(\log(n/\varepsilon))$, and runs in time $t\cdot\operatorname{polylog}(\Delta_{t})\mathopen{}\mathclose{{\left(\ell_{1}+\log^{2}(t/\varepsilon)}}\right)=O(n)$, assuming $c$ is small enough. Again, $Z_{2}$ is $\varepsilon^{\prime}$-close to an exact block source for every fixing of the seed.
For our next and final step, of performing the block-source extraction itself, set $d_{3}=c_{\mathsf{E}}\log(\ell_{t}/\varepsilon_{\mathsf{Ext}})$ where $c_{\mathsf{E}}$ is the constant guaranteed by lemma˜2.13. Also, let $\varepsilon_{\mathsf{Ext}}=\frac{\varepsilon^{\prime}}{6t}$, and $\theta$ will be a constant whose value will be later determined. We will use the following extractors:
  • Let $\mathsf{Ext}_{t}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\ell_{t}}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{3}}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m_{t}=(1+\theta)d_{3}}$ be the $(k_{t}=\delta^{\prime}\ell_{t},\varepsilon_{\mathsf{Ext}})$ extractor guaranteed to us by lemma˜2.13. Notice that we need to satisfy $k_{t}\geq\theta d_{3}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})$. This can be satisfied making sure that $\varepsilon$ is at most $2^{-\Omega(\ell_{t})}$, where the hidden constant depends on $c_{\mathsf{E}}$.
@@
    be the $(k_{i}=\delta^{\prime}\ell_{i},\varepsilon_{\mathsf{Ext}})$ extractor guaranteed to us by lemma˜2.13, where $m_{i}=(1+\theta)m_{i+1}$. We need to make sure that $m_{i+1}\geq c_{\mathsf{E}}\log(\ell_{i}/\varepsilon_{\mathsf{Ext}})$ and that $k_{i}\geq\theta m_{i+1}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})$. To see that the latter holds, note that $k_{i}=\delta^{\prime}\cdot\ell_{1}\alpha^{i-1}\geq n^{\gamma/2}$ and that $\theta m_{i+1}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})=\theta(1+\theta)^{t-i}d_{3}+c_{\mathsf{E}}\log(1/\varepsilon_{\mathsf{Ext}})<n^{\gamma/2},$ if we choose $\theta$ to be a small enough constant (with respect to the constant $\frac{\log n}{t}$) and $\varepsilon$ to be, again, at most $2^{-\Omega(\ell_{t})}$.
Everything is in place to apply our block source extraction, lemma˜2.22, on $Z_{2}$ and an independent and uniform seed of length $d_{3}$. We get that $\mathsf{BExt}$ outputs $Z_{3}$ of length $m_{1}=n^{\Omega(1)}$, which is $2t\varepsilon_{\mathsf{Ext}}\leq\varepsilon^{\prime}$ close to uniform, and runs in time $O\mathopen{}\mathclose{{\left(\sum_{i=1}^{t}\ell_{i}\log\ell_{i}}}\right)=O(n).$ Recall that indeed, as lemma˜2.22 requires, all the $\mathsf{Ext}_{i}$-s output their seed.
@@ 4.1.4 Improving the output length
The extractor $\mathsf{Ext}_{\mathsf{short}}$ from lemma˜4.4 only outputs $n^{\Omega(1)}$ bits. Here, we will use an extractor $\mathsf{Ext}_{\mathsf{out}}$ that outputs a linear fraction of the entropy but requires a (relatively) long seed, and use lemma˜2.25 to boost the output length. For $\mathsf{Ext}_{\mathsf{out}}$, we will again use a sample-then-extract extractor, however this time, we can use independent samples to create a block source with exponentially decreasing blocks. This setting is easier, and we can simply use the original [NZ96] construction. Since a similar construction will be analyzed later in the paper (including a time complexity analysis), we choose to employ it instead of revisiting [NZ96]. We state it formally as a corollary below.
@@
There exists a constant $C\geq 1$ such that for any constants $\tau,c\in(0,1)$, any large enough positive integer $n$ and any $\varepsilon\geq 2^{-n^{c}}$, there exists a strong $(k=(1-\tau)n,\varepsilon)$ extractor
@@
where $d=O(\log n\cdot\log(n/\varepsilon))$, and $m=k/C$. Moreover, given inputs $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$, we can compute $\mathsf{Ext}_{\mathsf{out}}(x,y)$ in time $\widetilde{O}(n)$.
The correctness of corollary˜4.5 follows from corollary˜4.10 applied with $i=1$ (which is indeed non-recursive), without the need for a preliminary condensing step. [^26] As mentioned earlier, corollary 4.5 can already be deduced from [NZ96] (modulo the tight runtime analysis), with a slightly worse seed of $d=O(\log^{2}n\cdot\log(n/\varepsilon))$, which would not change the parameters of our overall construction. There, they first convert $X$ into a block source $Z$ using $\log n$ independent samples from a $k$-wise independent sample space, for $k\approx\log(n/\varepsilon)$. The block source $Z$ has decreasing blocks, so the standard block source extraction scheme can then be employed. The fact that this procedure can be implemented in time $\widetilde{O}(n)$ follows easily from the runtime analysis of primitives in our work (specifically, item 1 of lemma 2.1, and lemma 2.13).
Plugging-in $\mathsf{Ext}_{\mathsf{out}}$ and $\mathsf{Ext}_{\mathsf{short}}$ into lemma˜2.25 readily gives the following result.
@@
There exist constants $\tau,c\in(0,1)$ such that for every positive integer $n$, and any $2^{-n^{c}}\leq\varepsilon\leq\frac{1}{n}$, there exists a $(k=(1-\tau)n,\varepsilon)$ extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ where $d=O(\log(n/\varepsilon))$, and $m=ck$. Moreover, given inputs $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$, we can compute $\mathsf{Ext}(x,y)$ in time $\widetilde{O}(n)$.
To boost the output length in lemma˜4.6 from $\Omega(k)$ to $(1-\eta)k$ for any constant $\eta>0$, we apply lemma˜2.21 a constant number of times depending only on $\eta$ (that is, we simply apply $\mathsf{Ext}$ with independent seeds and concatenate the outputs). To then go from any min-entropy requirement $k$ to entropy rate $1-\tau$, we first apply the KT condenser from theorem˜3.14. Since $\varepsilon\geq 2^{-k^{c}}$ we also have that $\varepsilon\geq 2^{-n^{0.1}}$ if $c<0.1$, and so the KT condenser does not require preprocessing. Furthermore, we can ensure that $k\geq C^{\prime}\log^{2}(n/\varepsilon)$ with $C^{\prime}>0$ a sufficiently large constant so that the conditions for applying the KT condenser are satisfied whenever $n$ is larger than some constant.
This finally gives us our main theorem for this section, theorem˜4.1, apart from the strongness property, which we now discuss.
@@
In what follows, we refer to the itemized list in the beginning of the section. The condensing step, item˜1, is strong, since we use strong condensers. Next, the block source creators of lemmas˜4.2 and 4.3 are strong, so items˜2 and 3 hold in a strong manner as well. item˜4 readily gives a strong extractor. For the output-extending phase, lemma˜2.25 tells us that the extractor from lemma˜4.6 is strong. Finally, we apply that extractor several times with independent seeds, and the strongness of that procedure is guaranteed from lemma˜2.21.